OpenStudy (anonymous):

f(x) = x^2 + x - 30 what are the zeros of the equation?

OpenStudy (anonymous):

@iambatman I don't need the answer, I just need someone to walk me through this step by step cause I know how to get the answer however I don't know how to get there

OpenStudy (anonymous):

I know I have to use the quadratic formual. but I don't really understand how to set it up correctly

OpenStudy (danjs):

The zeros are where the thing intersects the x-axis. The place(s) where y=f(x) = 0 set f(x) = 0 0 = x^2 + x - 30

OpenStudy (anonymous):

oh okay.

OpenStudy (danjs):

the general form for a quadratic is Ax^2 + Bx + C, where A, B and C are the coefficients on each of the terms, the last being Cx^0

OpenStudy (danjs):

That one you have can be factored pretty easy if you dont want to use the quadratic formula. 0 = (x + )( x - )

OpenStudy (danjs):

a*b = 30 and a - b = +1 What factors of 30, add together to give +1

OpenStudy (anonymous):

I suppose I'm still a bit lost about what exactly to do after I set f(x) to 0

OpenStudy (anonymous):

I could easily start guessing at what x is but that doesn't seem right

OpenStudy (danjs):

30 = a*b = 5*6 +1 = a - b = 6 - 5

OpenStudy (anonymous):

do I just start plugging in values for x?

OpenStudy (danjs):

You can always use the quadratic formula if you want. This one is easy to factor because the zeroes are whole number integers. f(x) = (x+6)(x-5)

OpenStudy (danjs):

5*6 is 30 and 6-5 is 1

OpenStudy (danjs):

or you can use the formula in any case \[x = \frac{ -B \pm \sqrt{B^2 - 4*A*C} }{ 2A }\] A= 1 B= 1 C = -30

OpenStudy (anonymous):

so with the other one I'd say x * x^2 = -30?

OpenStudy (danjs):

The function is f(x) = x^2 + x - 30 because the coefficient on the x^2 is a 1, you can write...

OpenStudy (danjs):

\[f(X) = (x \pm a)(x \pm b)\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

then what do I do?

OpenStudy (anonymous):

OHHH I finally got what you were saying uptop

OpenStudy (danjs):

Then you write the factors of 30 down... 30 * 1 6 * 5 2*15 3*10 ------------ Think, which of those factors will add or subtract to give you the value of B, in this case B is +1

OpenStudy (danjs):

the factors 6*5 are the ones 6*5 = 30 and 6-5 = 1

OpenStudy (danjs):

x^2 +x - 30 = (x +6)(x-5)

OpenStudy (anonymous):

ohhh okay. that makes sense now :)

OpenStudy (danjs):

In any case though, you can use the quadratic formula and get the answer for the zeros... \[x = \frac{ -B \pm \sqrt{B^2 - 4*A*C} }{ 2A } = \frac{ -1 \pm \sqrt{1 - (4*1*-30)} }{ 2*1 } = \frac{ -1 \pm \sqrt{121} }{ 2 } = \frac{ -1 \pm 11 }{ 2 }\]

OpenStudy (danjs):

\[x=\frac{-1+11 }{ 2 }~~~or~~~x=\frac{ -1-11 }{ 2 }\] x=5 or x=-6

OpenStudy (anonymous):

okay thank you for you help DanJ! :)

OpenStudy (danjs):

Which tells you the factors (x - 5)(x+6) = 0

OpenStudy (danjs):

(x-5) = 0 or (x+6) = 0 x=5 or x=-6

OpenStudy (danjs):

you're welcome, anytime